### v1

# Infinite log-concavity and higher order Turán inequality for Speyer’s $g$-polynomial of uniform matroids

Abstract.
Let $U_{n,d}$ be the uniform matroid of rank $d$ on $n$ elements. Denote by $g_{U_{n,d}}(t)$ the Speyer’s $g$-polynomial of $U_{n,d}$. The Turán inequality and higher order Turán inequality are related to the Laguerre-Pólya ($\mathcal{L}$-$\mathcal{P}$) class of real entire functions, and the $\mathcal{L}$-$\mathcal{P}$ class has close relation with the Riemann hypothesis. The Turán type inequalities have received much attention. Infinite log-concavity is also a deep generalization of Turán inequality with different direction. In this paper, we mainly obtain the infinite log-concavity and the higher order Turán inequality of the sequence $\{g_{U_{n,d}}(t)\}_{d=1}^{n-1}$ for $t>0$. In order to prove these results, we show that the generating function of $g_{U_{n,d}}(t)$, denoted $h_{n}(x;t)$, has only real zeros for $t>0$. Consequently, for $t>0$, we also obtain the $\gamma$-positivity of the polynomial $h_{n}(x;t)$, the asymptotical normality of $g_{U_{n,d}}(t)$, and the Laguerre inequalities for $g_{U_{n,d}}(t)$ and $h_{n}(x;t)$.
Key words and phrases:
Speyer’s $g$-polynomial, uniform matroid, infinitely log-concavity, higher order Turán inequality, real zeros, interlacing property, asymptotic normality
1991 Mathematics Subject Classification:
Primary 05B35, 26C10, 33F10, 05A20, 60F05
1. Introduction
In the study of the tropical analogues of the notions of linear spaces and Plücker coordinates, Speyer [46] found that all constructible tropical linear spaces have the same $f$-vector and are series-parallel. He also conjectured that this $f$-vector is maximal for all tropical linear spaces, with equality precisely for the series-parallel tropical linear spaces. In the study of his $f$-vector conjecture, Speyer [47] discovered an invariant, denoted by $g_{M}(t)$, for a (loopless and coloopless) $\mathbb{C}$-realizable matroid $M$. By using this invariant, Speyer proved bounds on the complexity of Kapranov’s Lie complexes [32], Hacking, Keel and Tevelev’s very stable pairs [29], and tropical linear spaces defined in [46]. It turns out that this invariant is of independent combinatorial interest. In a subsequent work of Fink and Speyer [24], the invariant $g_{M}(t)$ was defined for an arbitrary (loopless and coloopless) matroid $M$. López de Medrano, Rincón and Shaw [37] conjectured a Chow-theoretic formula for Speyer’s $g$-polynomial, which was later proved by Berget, Eur, Spink and Tseng [3]. Ferroni [22] provided a combinatorial way of computing $g_{M}(t)$ for an arbitrary Schubert matroid $M$. Recently, Ferroni and Schröter [23] proved that the coefficients of $g_{M}(t)$ are nonnegative for any sparse paving matroid $M$.
It is worthwhile to mention that the $g$-polynomial of uniform matroids has been computed by Speyer [47], and this polynomial plays a key role in his proof of the $f$-vector conjecture in the case of a tropical linear space realizable in characteristic zero. Denote by $U_{n,d}$ the uniform matroid of rank $d$ on $n$ elements. Speyer [47] obtained the following result.
Proposition 1.1 ([47, Proposition 10.1]).
For any $n\geq 2$ and $11\leq d\leq n-$ 1, the $g$-polynomial of the uniform matroid $U_{n,d}$ is given by
(1.1) $g_{U_{n,d}}(t)=\sum_{i=1}^{\min(d,n-d)}\frac{(n-i-1)!}{(d-i)!(n-d-i)!(i-1)!}\,t^{i}.$
Recently, Zhang and Zhao [58] showed the real-rootedness of the polynomial $g_{U_{n,d}}(t)$ and asymptotic normality of the coefficient of $g_{U_{n,[n/2]}}(t)$. This motivated us to explore more properties of this polynomial.
The objective of this paper is to prove the infinite log-concavity and higher order Turán inequality of Speyer’s $g$-polynomial $g_{U_{n,d}}(t)$ for any given $t>0$.
A sequence $\{a_{n}\}_{n\geq 0}$ of real numbers is said to be log-concave if $a_{n}^{2}-a_{n-1}a_{n+1}\geq 0$ for $n\geq 1$. This inequality is also called the Turán inequality [51, 52] or the Newton’s inequality [14, 15, 40]. A polynomial is said to be log-concave if the sequence of its coefficients is log-concave, see Brenti [4] and Stanley [48].
For a sequence $\{a_{n}\}_{n\geq 0}$, define an operator $\mathfrak{L}$ by $\mathfrak{L}(\{a_{n}\}_{n\geq 0})=\{b_{n}\}_{n\geq 0}$, where $b_{n}=a_{n}^{2}-a_{n-1}a_{n+1}$ for $n\geq 0$, with the convention that $a_{-1}=0$. A sequence $\{a_{n}\}_{n\geq 0}$ is said to be $k$-log-concave if the sequence $\mathfrak{L}^{j}\left(\{a_{n}\}_{n\geq 0}\right)$ is nonnegative for each $1\leq j\leq k$, and $\{a_{n}\}_{n\geq 0}$ is said to be $\infty$-log-concave if $\mathfrak{L}^{k}\left(\{a_{n}\}_{n\geq 0}\right)$ is nonnegative for any $k\geq 1$.
The notion of infinite log-concavity was introduced by Boros and Moll [6]. They also conjectured that a sequence arose from their study of evaluation of a quartic integral, named $\{d_{\ell}(m)\}_{\ell=0}^{m}$, is infinitely log-concave, which is still open. This conjecture has inspired many interesting works. For example, Brändén [9] proved the following result and hence confirmed a conjecture independently posed by Fisk [25], McNamara and Sagan [39] and Stanley; see [9].
Theorem 1.2.
[9] If a polynomial $\sum_{k=0}^{n}a_{k}x^{k}$ has only real and negative zeros, then so does the polynomial $\sum_{k=0}^{n}(a_{k}^{2}-a_{k-1}a_{k+1})z^{k}$ with the convention that $a_{-1}=a_{n+1}=0$. In particular, the sequence $\{a_{k}\}_{k=0}^{n}$ is infinitely log-concave.
The first main result of this paper is as follows.
Theorem 1.3.
Given $t>0$, the sequence $\{g_{U_{n,d}}(t)\}_{d=1}^{n-1}$ is infinitely log-concave for $n\geq 4$.
A real sequence $\{a_{n}\}_{n\geq 0}$ is said to satisfy the higher order Turán inequality or cubic Newton inequality if for $n\geq 1$,
(1.2) $4(a_{n}^{2}-a_{n-1}a_{n+1})(a_{n+1}^{2}-a_{n}a_{n+2})-(a_{n}a_{n+1}-a_{n-1}a_{n+2})^{2}\geq 0,$
see [18, 40, 44]. The Turán inequality and the higher order Turán inequality are related to the Laguerre-Pólya class of real entire functions, see [18, 51]. A real entire function $\psi(x)=\sum_{n=0}^{\infty}a_{n}\frac{x^{n}}{n!}$ is said to belong to the Laguerre-Pólya class, denoted by $\psi\in\mathcal{L}$-$\mathcal{P}$, if $\psi(x)=cx^{m}e^{-\alpha x^{2}+\beta x}\prod_{k=1}^{\infty}(1+x/x_{k})e^{-x/x_{k}},$ where $c,\beta,x_{k}$ are real, $\alpha\geq 0$, $m$ is a nonnegative integer and $\sum x_{k}^{-2}<\infty$. Jensen [31] proved that a real entire function $\psi(x)\in\mathcal{L}$-$\mathcal{P}$ if and only if for any integer $n>0$, the $n$-th associated Jenson polynomial $J_{n}(x)=\sum_{k=0}^{n}\binom{n}{k}a_{k}x^{k}$ are hyperbolic, i.e., $J_{n}(x)$ has only real zeros. This result was also obtained by Pólya and Schur [43]. Besides, if a real entire function $\psi\in\mathcal{L}$-$\mathcal{P}$, then its Maclaurin coefficients are log-concave, see [14, 16]. Moreover, Dimitrov [18] proved that the higher order Turán inequality (1.2), an extension of Turán inequality, is also a necessary condition for $\psi\in\mathcal{L}{\textrm{-}}\mathcal{P}$.
The $\mathcal{L}$-$\mathcal{P}$ class has close relation with the Riemann hypothesis. Let $\zeta$ and $\Gamma$ denote the Riemann zeta function and the gamma function, respectively. The Riemann $\xi$-function is defined by $\xi(iz)=\frac{1}{2}\left(z^{2}-\frac{1}{4}\right)\pi^{-\frac{z}{2}-\frac{1}{4}}\Gamma\left(\frac{z}{2}+\frac{1}{4}\right)\zeta\left(z+\frac{1}{2}\right)$, see Boas [10]. It is well known that the Riemann $\xi$-function is an entire function of order $1$ and can be restated as
(1.3) $\frac{1}{8}\,\xi\!\left(\frac{x}{2}\right)=\sum\limits_{k=0}^{\infty}(-1)^{k}\,\hat{b}_{k}\,\frac{x^{2k}}{(2k)!},$
[…]
$\xi_{1}(z)=\sum_{k=0}^{\infty}\frac{k!}{(2k)!}\hat{b}_{k}\frac{z^{k}}{k!}.$
So the Riemann hypothesis holds if and only if $\xi_{1}(z)\in\mathcal{L}$-$\mathcal{P}$. See [12, 15, 18] for more details.
For the deep relation stated above, the higher order Turán inequality has received great attention. Many interesting sequences were showed to satisfy the higher order Turán inequality (1.2). For instance, the Riemann $\xi$-function was proved to satisfy (1.2) by Dimitrov and Lucas [19]. By using the Hardy-Ramanujan-Rademacher formula, Chen, Jia and Wang [12] proved that the partition function $p(n)$ satisfies the higher order Turán inequality for $n\geq 95$ and the $3$-rd associated Jensen polynomials $\sum_{k=0}^{3}\binom{3}{k}p(n+k)x^{k}$ have three distinct zeros, and hence confirm a conjecture of Chen [11]. Griffin, Ono, Rolen and Zagier [28] showed that Jensen polynomials for a large family of functions, including those associated to $\zeta(s)$ and the partition function $p(n)$, are hyperbolic for sufficiently large $n$, which gave evidence for Riemann hypothesis. For more results on this topic, please see [55, 53, 54, 20] and the literature cited therein.
The second main result of this section is as follows.
Theorem 1.4.
Given $t>0$, the sequence $\{g_{U_{n,d}}(t)\}_{d=1}^{n-1}$ satisfies the higher order Turán inequality (1.2) for $n\geq 5$.
In order to prove Theorems 1.3 and 1.4, we investigate the real-rootedness of the following generating function of Speyer’s $g$-polynomial
(1.4) $h_{n}(x;t):=\sum_{d=1}^{n-1}g_{U_{n,d}}(t)x^{d},$
where $t$ is treated as a nonzero parameter. Clearly, $h_{n}(x;t)$ is a polynomial in $x$ with leading term $tx^{n-1}$. If $h_{n}(x;t)$ has only real zeros, then we can derive Theorem 1.3 from Theorem 1.2 of Brändén. As will be seen, the real-rootedness of $h_{n}(x;t)$ also plays a key role in the proof of Theorem 1.4.
For notational convenience, let $g_{n,d}(t)=g_{U_{n,d}}(t)$ and $S_{i}(n,d)=[t^{i}]g_{n,d}(t)$, the coefficient of $t^{i}$ in $g_{n,d}(t)$. From (1.1) it immediately follows that $g_{n,d}(t)=g_{n,n-d}(t)$. So the polynomial $h_{n}(x;t)$ is symmetric for any $n\geq 3$. Clearly, $S_{i}(n,d)=\frac{(n-i-1)!}{(d-i)!(n-d-i)!(i-1)!}=\binom{n-i-1}{d-i}\binom{n-d-1}{i-1}.$ Since the binomial coefficient $\binom{n}{k}=0$ for $k>n$, we can rewrite (1.4) as
(1.5) $h_{n}(x;t):=\sum_{d=1}^{n-1}\sum_{i=1}^{d}\binom{n-i-1}{d-i}\binom{n-d-1}{i-1}\,t^{i}x^{d}.$
The remainder of this paper is organized as follows. In Section 2, we first obtain several recurrence relations of $h_{n}(x;t)$ by using a computer algebraic system developed by Koutschan [33]. With one of these recurrences and a result due to Liu and Wang [36], we show that the polynomial sequence $\{h_{n}(x;t)\}_{n\geq 2}$ has interlacing zeros for $t>0$, in Section 3, and hence obtain the real-rootedness of $h_{n}(x;t)$ for $t>0$ and $n\geq 2$. In Section 4, we complete the proofs of Theorems 1.3 and 1.4. The number $g_{n,d}(1)$ possesses specific geometric interpretations which will also be stated in detail in Section 5. It turns out that the polynomial $h_{n}(x;t)$ enjoys more properties. In Sectin 6, we prove the $\gamma$-positivity of $h_{n}(x;t)$. In Section 7, we show that the coefficient of $h_{n}(x;t)$, namely the Speyer’s $g$-polynomial $g_{n,d}(t)$, is asymptotically normal by local and central limit theorems based on a criterion given by Bender [2] and Harper [30]. Finally, in Section 8, we give Laguerre inequalities related to $g_{n,d}(t)$ and $h_{n}(x;t)$ and conclude this paper with a conjecture.
2. Recurrence Relations
The aim of this section is to show some recurrence relations for the generating function of Speyer’s $g$-polynomial $h_{n}(x;t)$. As will be seen, these recurrences play an important role in the proofs of the real-rootedness, the $\gamma$-positivity, and the asymptotic normality presented in Sections 3, 6, and 7, respectively. And the real-rootedness of $h_{n}(x;t)$ will lead to Theorems 1.3 and 1.4. The main result of this section is as follows.
Theorem 2.1.
Let $h_{n}(x;t)$ be a function of $x$ defined by (1.5). Then for $n\geq 2$, we have
(2.1) $h_{n+2}(x;t)$ $\,=(1+x)h_{n+1}(x;t)+txh_{n}(x;t),$
(2.2) $A\,h_{n+1}(x;t)$ $\,=B\,h_{n}(x;t)+C\,h_{n}^{\prime}(x;t),$
(2.3) $\widetilde{A}\,h_{n+1}(x;t)$ $\,=\widetilde{B}\,h_{n}(x;t)+C^{2}\,h_{n}^{\prime\prime}(x;t),$
where
$A=$ $\,(n-1)(x-1),$
$B=$ $\,-n(2tx+x+1),$
$C=$ $\,x(x^{2}+4tx+2x+1),$
and
$\widetilde{A}=$ $\,(n-1)\big{(}(n-2)x^{3}+(4nt+n-2t+2)x^{2}-(4nt+n-6t-4)x-n\big{)},$
$\widetilde{B}=$ $\,-n\big{(}(nt-5t-2)x^{3}+(4nt^{2}+2nt-8t^{2}+n-6t-5)x^{2}$
$\qquad\qquad\qquad+(5nt+2n-9t-4)x+n-1\big{)}.$
Proof.
Since the coefficients of the recurrence (2.3) looks a little complicated, a symbolic proof based on a computer algebra system will be more concise. We shall apply the package HolonomicFunctions [^1] See https://www3.risc.jku.at/research/combinat/software/ergosum/RISC/HolonomicFunctions.html. for Mathematica developed by Koutschan [33] to prove Theorem 2.1. Let us first load this package.
In[1]:=$<<\mathrm{RISC^{\prime}HolonomicFunctions^{\prime}}$
HolonomicFunctions Package version 1.7.3 (21-Mar-2017)
written by Christoph Koutschan
Copyright Research Institute for Symbolic Computation (RISC),
Johannes Kepler University, Linz, Austria
Then run the command Annihilator[$expr$, $ops$] which computes annihilating relations for $expr$ with respect to the Ore operator(s) $ops$. In the following input, $expr=h_{n}(x;t)$ and $ops$ consists of S$[n]$ and Der$[x]$, where S$[n]$ stands for the forward shift operator such that S$[n]h_{n}(x;t)=h_{n+1}(x;t)$ and Der$[x]$ represents the operator partial derivative with respect to $x$.
In[2]:=$\mathrm{ann=Annihilator}[\mathrm{Sum}[\mathrm{Sum}[\mathrm{Binomial}[n-i-1,d-i]\mathrm{Binomial}[n-d-1,i-1]t^{i},\hfil\break\hbox{}\qquad\{i,1,d\}]x^{d},\{d,1,n-1\}],\{\mathrm{S}[n],\mathrm{Der}[x]\}]$
Out[2]=$\{(1-n-x+nx)S_{n}+(-x-2x^{2}-4tx^{2}-x^{3})D_{x}+(n+nx+2ntx),\hfil\break\hbox{}(x^{2}+x^{3}+4tx^{3}-x^{4}-4tx^{4}-x^{5})D_{x}^{2}+(-nx+4x^{2}-nx^{2}+6tx^{2}-4ntx^{2}+2x^{3}+nx^{3}-2tx^{3}\hfil\break\hbox{}+4ntx^{3}-2x^{4}+nx^{4})D_{x}+(n-3nx+n^{2}x-ntx+n^{2}tx-n^{2}x^{2}-ntx^{2}-n^{2}tx^{2})\}$
The output Out[2] gives a list of Ore Polynomial expressions which form a Gröbner basis of an annihilating left ideal for $h_{n}(x;t)$, where $S_{n}={\rm S}[n]$ and $D_{\!x}={\rm Der}[x]$. For more information of the Ore algebra and Ore polynomial, please see Ore [41] and Koutschan [33].
Next, we employ the command OreReduce[$opoly$, $ann$] which reduces the Ore polynomial $opoly$ modulo the Gröbner basis $ann$. In the following input, $opoly:={\rm S}[n]^{2}-(x+1){\rm S}[n]-tx$ which corresponds to the Ore polynomial of recurrence (2.1) and $ann$ is given by Out[2].
In[3]:=$\mathrm{OreReduce}[\mathrm{S}[n]^{2}-(x+1)\mathrm{S}[n]-tx,\mathrm{ann}]$
Out[3]=$0$
The return is zero, which means that $opoly$ is contained in the left ideal generated by the elements of $ann$, and hence $\big{(}{\rm S}[n]^{2}-(x+1){\rm S}[n]-tx\big{)}h_{n}(x;t)=0$, that is, the recurrence (2.1) holds.
To prove (2.2), set $opoly:=(n-1)(x-1){\rm S}[n]-x(x^{2}+4tx+2x+1){\rm Der}[x]+n(2tx+x+1)$ and run the following command:
In[4]:=$\mathrm{OreReduce}[(n-1)(x-1)\mathrm{S}[n]-x(x^{2}+4tx+2x+1)\mathrm{Der}[x]+n(2tx+x+1),\mathrm{ann}]$
Out[4]=$0$
This means that (2.2) is valid.
Finally, by setting $opoly=\widetilde{A}\,{\rm S}[n]-\widetilde{B}-C^{2}{\rm Der}[x]^{2}$ where $\widetilde{A}$, $\widetilde{B}$, and $C$ are given in Theorem 2.1, we have
In[5]:=$\mathrm{OreReduce}[(n-1)((n-2)x^{3}+(4nt+n-2t+2)x^{2}-(4nt+n-6t-4)x\hfil\break\hbox{}\qquad-n)\mathrm{S}[n]+n((nt-5t-2)x^{3}+(4nt^{2}+2nt-8t^{2}+n-6t-5)x^{2}\hfil\break\hbox{}\qquad+(5nt+2n-9t-4)x+n-1)-x^{2}(x^{2}+4tx+2x+1)^{2}\mathrm{Der}[x]^{2},\mathrm{ann}]$
Out[5]=$0$
The Out[2] means that the recurrence (2.3) is true. This completes the proof. ∎
3. Real-rootedness and interlacing property
The objective of this section is to show the real-rootedness of the generating function of Speyer’s $g$-polynomial $h_{n}(x;t)$ for any $t>0$.
Theorem 3.1.
For any $t>0$, the polynomial $h_{n}(x;t)$ in $x$ has only real zeros for all $n\geq 2$.
Before proving Theorem 3.1, let us first recall some definitions and results on Sturm sequences. Let PF be the set of real-rooted polynomials with nonnegative coefficients, including any nonnegative constant for convenience. Given two polynomials $f(x),\,g(x)\in\mathrm{PF}$, assume $f(u_{i})=0$ and $g(v_{j})=0$. We say that $g(x)$ interlaces $f(x)$, denoted $g(x)\preceq f(x)$, if either $\deg f(x)=\deg g(x)=n$ and
(3.1) ${v_{n}\leq u_{n}\leq\cdots\leq v_{2}\leq u_{2}\leq v_{1}\leq u_{1}},$
or $\deg f(x)=\deg g(x)+1=n$ and
(3.2) ${u_{n}\leq v_{n-1}\leq u_{n-1}\leq\cdots\leq v_{2}\leq u_{2}\leq v_{1}\leq u_{1}}.$
We say that $g(x)$ interlaces $f(x)$ strictly, denoted $g(x)\prec f(x)$, if the inequalities in (3.1) or (3.2) hold strictly. Following Liu and Wang [36], we also let $a\preceq bx+c$ for any nonnegative $a,b,c$, and let $0\preceq f$ and $f\preceq 0$ for any $f\in\mathrm{PF}$. A polynomial sequence $\{f_{n}(x)\}_{n\geq 0}$ with each $f_{n}(x)\in\mathrm{PF}$ is said to be a generalized Sturm sequence if
${f_{0}(x)\preceq f_{1}(x)\preceq\cdots\preceq f_{n-1}(x)\preceq f_{n}(x)\preceq\cdots.}$
Liu and Wang [36] established the following sufficient condition for determining whether two polynomials have interlacing zeros.
Theorem 3.2.
[36, Theorem 2.3] Let $F(x),f(x),g_{1}(x),\ldots,g_{k}(x)$ be real polynomials satisfying the following conditions.
  • $(i)$
    There exist some real polynomials $\phi(x),\psi_{1}(x),\ldots,\psi_{k}(x)$ such that
    (3.3) $F(x)=\phi(x)f(x)+\psi_{1}(x)g_{1}(x)+\cdots+\psi_{k}(x)g_{k}(x),$
    and $\deg F(x)=\deg f(x)$ or $\deg F(x)=\deg f(x)+1$.
  • $(ii)$
    $f(x),\,g_{j}(x)$ are polynomials with only real zeros, and $g_{j}(x)\preceq f(x)$ for all $j$.
  • $(iii)$
    The leading coefficients of $F(x),g_{1}(x),\ldots,g_{k}(x)$ have the same sign.
Suppose that $\psi_{j}(r)\leq 0$ for each $j$ and each zero $r$ of $f(x)$. Then $F(x)$ has only real zeros and $f(x)\preceq F(x)$. In particular, if for each zero $r$ of $f(x)$, there exists an index $j$ such that $g_{j}(x)\prec f(x)$ and $\psi_{j}(r)<0$, then $f(x)\prec F(x)$.
We are now in a position to present the interlacing property of $h_{n}(x;t)$, which is stronger than Theorem 3.1.
Theorem 3.3.
For any $t>0$ the sequence $\{h_{n}(x;t)\}_{n\geq 2}$ is a generalized Sturm sequence.
Proof.
Fix $t>0$ and $n\geq 2$ throughout this proof. Denote $h_{n}(x):=h_{n}(x;t)$ for brief. We shall prove the desired result by using induction on $n$. First, by (1.5) we have $h_{2}(x)=tx$ and $h_{3}(x)=tx^{2}+tx$. Clearly, both $h_{2}(x)$ and $h_{3}(x)$ are real-rooted polynomials in $x$, and
$h_{2}(x)\preceq h_{3}(x).$
Assume that $h_{n}(x)$ and $h_{n+1}(x)$ are real-rooted polynomials and $h_{n}(x)\preceq h_{n+1}(x)$. It is sufficient to show
(3.4) $h_{n+1}(x)\preceq h_{n+2}(x).$
To this end, we shall employ Theorem 3.2. First, by recursion (2.1), we have
$h_{n+2}(x)$ $\,=(1+x)h_{n+1}(x)+txh_{n}(x)$
which is of the form of (3.3) with $k=1$, where
$F(x)=h_{n+2}(x),\quad f(x)=h_{n+1}(x),\quad g_{1}(x)=h_{n}(x),$
and
$\phi(x)=1+x,\quad\psi_{1}(x)=tx.$
By (1.5), we have that $\deg F(x)=\deg f(x)+1=n+1$. Thus, condition (i) of Theorem 3.2 is satisfied. By inductive hypothesis that $h_{n}(x)\preceq h_{n+1}(x)$, we have condition (ii) of Theorem 3.2 holds, too. Moreover, the leading coefficients of $F(x)$ and $g_{1}(x)$ are all positive by (1.5). Therefore, condition (iii) of Theorem 3.2 is also satisfied. It remains to show that $\psi_{1}(r)\leq 0$ for each zero $r$ of $f(x)$. Note that $f(x)$ has only nonpositive zeros since its coefficients are all nonnegative. Clearly, for $x\leq 0$, we have $\psi_{1}(x)\leq 0$.
It follows from Theorem 3.2 that (3.4) holds and $h_{n+2}(x)$ has only real zeros. This completes the proof. ∎
4. Proofs of main results
In this section, we complete the proofs of Theorems 1.3 and 1.4.
Proof of Theorem 1.3.
It is well-known that a real-rooted polynomial with nonnegative coefficients is log-concave. By Theorems 3.1, the polynomial $h_{n}(x;t)$ in $x$ has only real zeros for any $t>0$ and $n\geq 2$. It follows from Theorem 1.2 that the sequence $\{g_{U_{n,d}}(t)\}_{d=1}^{n-1}$ is infinitely log-concave for $n\geq 4$ and $t>0$. ∎
Mařík [38] showed that if $f(x)=\sum_{k=0}^{n}\frac{a_{k}}{k!(n-k)!}x^{k}$ has only real zeros for $n\geq 3$, then the sequence $\{a_{k}\}_{k=0}^{n}$ satisfies the higher order Turán inequality (1.2). Recently, Wang and Zhang [54] extended Mařík’s result and established the following sufficient conditions for proving (1.2) for nonnegative sequences.
Theorem 4.1.
[54, Theorem 1.2] Let $n\geq 3$ and $\{a_{k}\}_{k=0}^{n}$ be a sequence of positive integers. Let $f_{1}(x)=\sum_{k=0}^{n}a_{k}x^{k}$, $f_{2}(x)=\sum_{k=0}^{n}\frac{a_{k}}{k!}x^{k}$, and $f_{3}(x)=\sum_{k=0}^{n}\frac{a_{k}}{(n-k)!}x^{k}$. If one of $f_{i}(x)$, $1\leq i\leq 3$ has only real zeros, then the sequence $\{a_{k}\}_{k=0}^{n}$ satisfies the higher order Turán inequality (1.2).
We are now ready to show a proof of Theorem 1.4.
Proof of Theorem 1.4.
Note that Theorem 4.1 still holds for $a_{k}\in\mathbb{R}_{\geq 0}$, which can be verified by the argument in the proof of [54, Theorem 1.2]. By Theorems 3.1, we have the real-rootedness of the polynomial $h_{n}(x;t)$ for any $t>0$ and $n\geq 2$. Set $a_{k}:=g_{n,k}(t)$ in Theorem 4.1 with the convention that $a_{0}=a_{n}=0$. Then for any given $t>0$, we get the desired result that the sequence $\{g_{n,k}(t)\}_{k=1}^{n-1}$ satisfies the higher order Turán inequality (1.2) for $n\geq 5$. This completes the proof. ∎
5. Inequalities for $g_{n,d}(1)$
As noted by Speyer [46, 47], the number $S_{i}(n,d)=\binom{n-i-1}{d-i}\binom{n-d-1}{i-1}$ has interesting geometric interpretation. Therefore, $g_{n,d}(1)=\sum_{i=1}^{d}S_{i}(n,d)$ deserves more attention.
Let us recall some definitions and results showed by Speyer [46, 47]. Let $G(d,n)$ denote the Grassmannian of $d$-subspaces in $\mathbb{C}^{n}$. Let $K=\bigcup_{n=1}^{\infty}\mathbb{C}((t^{1/n}))$ be the filed of Puiseux series, and let $v:K^{*}\rightarrow\mathbb{Q}$ be the map such that if $x=\sum_{i\geq M}a_{i}t^{i/N}$ with $a_{M}\neq 0$ then $v(x)=M/N$. Assume that there exists a $K$-valued point of $G(d,n)$ none of whose Plücker coordinate $p_{I}(x)$ are zero. To helps to understand the combinatorial meaning of being a tropical Plücker vector, Speyer provided a simple polyhedral construction. Let $\Delta(d,n)$ denote the $(d,n)$-hypersimplex, the convex hull of the $\binom{n}{d}$ points $e_{i_{1}}+\cdots+e_{i_{d}}$ where $(i_{1},\ldots,i_{d})$ runs over $\binom{[n]}{d}$, with the convention that $[n]=\{1,2,\ldots,n\}$ and for any set $S$, $\binom{S}{d}$ is the set of $d$ element subsets of $S$. Let $I\mapsto P_{I}$ be a function $\binom{[n]}{d}\rightarrow\mathbb{Q}$. Speyer [47] defined a polyhedral subdivision of $\Delta(d,n)$ as follows: Let $Q$ be the convex hull in $\Delta(d,n)\times\mathbb{R}$ of the points $(e_{i_{1}}+\cdots+e_{i_{d}},P_{i_{1}\ldots i_{d}})$ and let $Q+\mathbb{R}_{\geq 0}$ denote the Minkowski sum of $Q$ with $\{0\}\times\mathbb{R}_{\geq 0}\subset\mathbb{R}^{n}\times\mathbb{R}$. Take the facets of $Q+\mathbb{R}_{\geq 0}$ whose outward pointing normal vectors have negative components in the last coordinate and project them down to $\Delta(d,n)$. This gives a polyhedral subdivision $\mathcal{D}_{P}$ of $\Delta(d,n)$ which is known as the regular subdivision associated to $P$. Then $P$ is a tropical Plücker vector if and only if, for each face $F$ of $\Delta$, the vertices of $F$ form the bases of a matroid. When this condition holds, the number of $i$-dimensional bounded faces of the corresponding tropical linear space is the number of $(n-i)$-dimensional interior faces in $\mathcal{D}$. Speyer obtained the following result and hence confirmed his $f$-vector conjecture [46] in the case of a tropical linear space realizable in characteristic zero.
Theorem 5.1.
[47] Suppose that $P$ is a tropical Plücker vector arising as $v(p_{I}(x))$ for some $x\in G(d,n)(K)$. Then $\mathcal{D}_{P}$ has at most $\binom{n-i-1}{d-i}\binom{n-d-1}{i-1}$ interior faces of dimension $n-i$, with equality if and only if all of the facets of $\mathcal{D}_{P}$ correspond to series-parallel matroids.
Theorem 5.1 implies that $g_{n,d}(1)$ enumerates all the interior faces of $\mathcal{D}_{P}$ with dimension runs over $\{n-d,n-d+1,\ldots,n-1\}$ in the case that all of the facets of $\mathcal{D}_{P}$ correspond to series-parallel matroids. As a consequence of Theorems 1.3 and 1.4, we obtain the main result of this section.
Theorem 5.2.
For $n\geq 4$, the sequence $\{g_{n,d}(1)\}_{d=1}^{n-1}$ is infinitely log-concave. And for $n\geq 5$, the sequence $\{g_{n,d}(1)\}_{d=1}^{n-1}$ satisfies the higher order Turán inequality (1.2).
Speyer [47] also gave an alternative description of Theorem 5.1. Let $T$ be the torus $(\mathbb{C}^{*})^{n}/\textrm{diag}(\mathbb{C}^{*})$ which acts on $G(d,n)$. Denote by $\overline{Tx}$ the closure of the torus orbit through $x$. Speyer constructed a certain subscheme $Y$ of $G(d,n)$, and mentioned that $Y$ is a union of torus orbit closures $Y=\bigcup_{j}\overline{Tx_{j}^{1}}$, one for each facet of $\mathcal{D}$. The $\overline{Tx_{j}^{1}}$ are glued along smaller torus orbit closures indexed by the smaller interior faces of $\mathcal{D}$. That is, the structure sheaf $\mathcal{O}_{Y}$ of $Y$ fits into an exact complex of sheaves
$0\rightarrow\mathcal{O}_{Y}\rightarrow\bigoplus_{j=1}^{f_{1}}\mathcal{O}_{\overline{Tx_{j}^{1}}}\rightarrow\cdots\rightarrow\bigoplus_{j=1}^{f_{i}}\mathcal{O}_{\overline{Tx_{j}^{i}}}\rightarrow\cdots\rightarrow\bigoplus_{j=1}^{f_{n}}\mathcal{O}_{\overline{Tx_{j}^{n}}}\rightarrow 0$
where each $x_{j}^{i}$ is a point of $G(d,n)$ such that the torus orbit $Tx_{j}^{i}$ is $(n-i)$-dimensional. Denote by $f_{i}$ the number of strata of dimension $n-i$. Then Theorem 5.1 reads
$f_{i}\leq\binom{n-i-1}{d-i}\binom{n-d-1}{i-1}$
with equality if and only if each $\overline{Tx_{k}^{1}}$ corresponds to a series-parallel matroid.
Clearly, one can see that if each $\overline{Tx_{k}^{1}}$ corresponds to a series-parallel matroid, then $g_{n,d}(1)=\sum_{i}f_{i}$ and the number $\sum_{i}f_{i}$ satisfies Theorem 5.2, too.
6. The gamma-positivity
In this section, we show the $\gamma$-positive of the polynomial $h_{n}(x;t)$. Given a polynomial $f(x)=a_{r_{1}}x^{r_{1}}+\cdots+a_{r_{s}}x^{r_{s}}\in\mathbb{R}[x]$ with $r_{1}\leq\cdots\leq r_{s}$ and $a_{r_{1}},a_{r_{s}}\neq 0$, the darga of $f(x)$ is defined to be ${\rm dar}(f):=r_{1}+r_{s}$, see Zeilberger [57]. The polynomial $f(x)$ is called palindromic of darga $n$ if $r_{1}+r_{s}=n$ and $a_{r_{1}+k}=a_{r_{s}-k}$ for all $k$. Let $\mathcal{P}_{n}$ denote the palindromic polynomials of darga $n$ with the convention that ${\rm dar}(0)$ can be any nonnegative integer. Sun, Wang and Zhang [50, Theorem 2.8] showed that $\mathcal{P}_{n}$ forms a linear space of dimension $\lfloor n/2\rfloor+1$ and $\{x^{i}(1+x)^{n-2i}|i=0,1,\ldots,\lfloor n/2\rfloor\}$ is a basis of $\mathcal{P}_{n}$. Let $f(x)\in\mathcal{P}_{n}$ and
$f(x)=\sum_{i=0}^{\lfloor n/2\rfloor}\gamma_{i}x^{i}(1+x)^{n-2i}.$
We say that $f(x)$ is $\gamma$-positive of darga $n$ if all the $\gamma_{i}$ are nonnegative.
The notion of $\gamma$-positivity was first introduced by Fota and Schützenberger [26] in studying the classical Eulerian polynomials. The works of Stembridge [49], Brändén [7, 8] and Gal [27] made it clear that the concept of $\gamma$-positivity is of independently interest which provides a useful approach to the problem of unimodality for symmetric polynomials. As mentioned by Athanasiadis [1, P. 2], symmetry and real-rootedness of a polynomial $f(x)\in\mathbb{R}_{\geq 0}[x]$ implies its $\gamma$-positivity, see also [7, Lemma 4.1] and [27, Remark 3.1.1].
Clearly, the polynomial $h_{n}(x;t)=\sum_{d=1}^{n-1}g_{n,d}(t)x^{d}\in\mathcal{P}_{n}$ for any $t\in\mathbb{R}$, that is, $g_{n,1+k}(t)=g_{n,n-1-k}(t)$ for $10\leq k\leq n-$ 1 and $t\in\mathbb{R}$. It follows from Theorem 3.1 and the above statement that $h_{n}(x;t)$ is $\gamma$-positive for $t>0$. More precisely, we have the following result.
Theorem 6.1.
Let $t>0$. For $n\geq 2$, we have
(6.1) $h_{n}(x;t)=\sum_{i=0}^{\lfloor n/2\rfloor}\gamma_{n,i}x^{i}(1+x)^{n-2i},$
where $\gamma_{n,i}$ satisfies the following recurrence relation for $n\geq 4$ and $0\leq i\leq\lfloor n/2\rfloor$,
(6.2) $\gamma_{n,i}=\gamma_{n-1,i}+t\,\gamma_{n-2,i-1}$
with initial conditions $\gamma_{2,0}=\gamma_{3,0}=0$, $\gamma_{2,1}=\gamma_{3,1}=t$ and the conventions that $\gamma_{n,-1}=\gamma_{n,\lfloor n/2\rfloor+1}=0$ for $n\geq 2$.
Proof.
We shall prove Theorem 6.1 by using induction on $n$. First by (1.4), we have $h_{2}(x;t)=tx$, $h_{3}(x;t)=tx(1+x)$ and $h_{4}(x;t)=tx(1+x)^{2}+t^{2}x^{2}$. Clearly, (6.1) holds for $n=2,3,4$, and $\gamma_{2,0}=\gamma_{3,0}=0$, $\gamma_{2,1}=\gamma_{3,1}=t$. Moreover, it is easy to verify that the recurrence (6.2) holds for $n=4$ and $0\leq i\leq 2$. Assume (6.1) and (6.2) hold true for some $n-1$ and $n$ where $n\geq 4$, we proceed to prove that
(6.3) $h_{n+1}(x;t)=\sum_{i=0}^{\lfloor(n+1)/2\rfloor}\big{(}\gamma_{n,i}+t\,\gamma_{n-1,i-1}\big{)}x^{i}(1+x)^{n+1-2i}.$
By the recurrence (2.1) given in Theorem 2.1, we have for some $n\geq 4$,
$h_{n+1}(x;t)$ $\,=(1+x)h_{n}(x;t)+txh_{n-1}(x;t).$
It follows from the inductive hypothesis that
$h_{n+1}(x;t)=$ $\,(1+x)\sum_{i=0}^{\lfloor n/2\rfloor}\gamma_{n,i}x^{i}(1+x)^{n-2i}+tx\sum_{i=0}^{\lfloor(n-1)/2\rfloor}\gamma_{n-1,i}x^{i}(1+x)^{n-1-2i}$
(6.4) $=$ $\,\sum_{i=0}^{\lfloor n/2\rfloor}\gamma_{n,i}x^{i}(1+x)^{n+1-2i}+t\sum_{i=1}^{\lfloor(n+1)/2\rfloor}\gamma_{n-1,i-1}x^{i}(1+x)^{n+1-2i}.$
Note that $\lfloor n/2\rfloor=\lfloor(n+1)/2\rfloor$ if $n$ is even and $\lfloor n/2\rfloor+1=\lfloor(n+1)/2\rfloor$ if $n$ is odd. In view of the convention that $\gamma_{n,-1}=\gamma_{n,\lfloor n/2\rfloor+1}=0$ for $n\geq 2$, we get (6.3) from (6.4). This completes the proof. ∎
Clearly, Theorem 6.1 implies that the polynomial $h_{n}(x;t)$ is $\gamma$-positivity of darga $n$ for any $t>0$ and $n\geq 2$. Besides, for $n\geq 2$,
$h_{n}(-1;t)=\left\{\begin{array}[]{ll}0,&\textrm{ if }n\textrm{ is odd},\\ (-1)^{n/2}\gamma_{n,n/2},&\textrm{ if }n\textrm{ is even}.\end{array}\right.$
This means that $(-1)^{n/2}h_{n}(-1;t)\geq 0$ for $t>0$ and $n\geq 2$.
Notice that $\gamma_{n,i}$ is a function of $t$. For $t=1$, it reduces to $\binom{n-i-1}{i-1}$ with $0\leq i\leq\lfloor n/2\rfloor$. The terms of $\{\binom{n-i-1}{i-1}\}_{i=1}^{\lfloor n/2\rfloor}$ correspond to the sequence A011973 and A169803 in the OEIS founded by Sloane [45].
7. Asymptotic normality
This section is devoted to the study of asymptotic normality of Speyer’s $g$-polynomial $g_{n,d}(t)$ for any given $t>0$. Theorem 3.1 implies that the sequence $\{g_{n,d}(t)\}_{d=1}^{n-1}$ is log-concave and unimodal for any $t>0$. The coefficients of many real-rooted polynomials appear to be asymptotically normal, see [2, 5, 13, 35, 58] for examples. Motivated by this fact, we investigate that whether the generating function of Speyer’s $g$-polynomials possess this asymptotic property.
Before showing the main result of this section, we first recall some concepts and results. Let $\{f_{n}(x)\}_{n\geq 0}$ be a sequence of polynomials with nonnegative coefficients where
(7.1) $f_{n}(x)=\sum_{k=0}^{n}a(n,k)x^{k}.$
The coefficient $a(n,k)$ is said to be asymptotically normal with mean $\mu_{n}$ and variance $\sigma_{n}^{2}$ by a central limit theorem if
[…]
where $p(n,k)=\frac{a(n,k)}{\sum_{j=0}^{n}a(n,j)}$. The coefficient $a(n,k)$ is said to be asymptotically normal with mean $\mu_{n}$ and variance $\sigma_{n}^{2}$ by a local limit theorem on the real set $\mathbb{R}$ if
$\lim\limits_{n\rightarrow\infty}\sup\limits_{x\in\mathbb{R}}\left|\sigma_{n}p(n,\lfloor\mu_{n}+x\sigma_{n}\rfloor)-\frac{1}{\sqrt{2\pi}}\exp({-{x^{2}}/{2}})\right|=0.$
The main result of this section is as follows.
Theorem 7.1.
The coefficients of $h_{n}(x;t)$, that is, Speyer’s $g$-polynomial $g_{n,d}(t)$, is asymptotically normal by local and central limit theorems for any $t>0$.
In order to prove Theorem 7.1, we shall employ the following criterion, see Bender [2] and Harper [30].
Theorem 7.2 ([2, Theorem 2]).
Let $\{f_{n}(x)\}_{n\geq 0}$ be a real-rooted polynomial sequence with nonnegative coefficients as in (7.1). Let
(7.2) $\mu_{n}=\frac{f_{n}^{\prime}(1)}{f_{n}(1)}\quad\textrm{and}\quad\sigma_{n}^{2}=\frac{f_{n}^{\prime\prime}(1)}{f_{n}(1)}+\mu_{n}-\mu_{n}^{2}.$
If $\sigma_{n}^{2}\rightarrow+\infty$ as $n\rightarrow+\infty$, then the coefficient of $f_{n}(x)$ is asymptotically normal with mean $\mu_{n}$ and variance $\sigma_{n}^{2}$ by local and central limit theorems.
We are now in a position to show a proof of Theorem 7.1.
Proof of Theorem 7.1.
Fix $t>0$ throughout this proof. For $n\geq 2$, let $f_{n}(x):=h_{n}(x;t)$, $\mu_{n}=\frac{f_{n}^{\prime}(1)}{f_{n}(1)}$ and $\sigma_{n}^{2}=\frac{f_{n}^{\prime\prime}(1)}{f_{n}(1)}+\mu_{n}-\mu_{n}^{2}$. By Theorem 3.1, we have $\{f_{n}(x)\}_{n\geq 2}$ is a real-rooted polynomial sequence. Clearly, the coefficients of $f_{n}(x)$ are nonnegative by (1.1). By Theorem 7.2 it is sufficient to show that $\sigma_{n}^{2}\rightarrow+\infty$ as $n\rightarrow+\infty$. To this end, we need first obtain the limitation of ${f_{n+1}(1)}/{f_{n}(1)}$. Substitute $x=1$ in the recurrence (2.1), we have
(7.3) $f_{n+2}(1)$ $\,=2f_{n+1}(1)+tf_{n}(1),\qquad n\geq 2.$
Let $r_{n}(x)=\frac{f_{n+1}(x)}{f_{n}(x)}$ for $n\geq 2$. Suppose that
$\lim_{n\rightarrow+\infty}r_{n}(1)=\lambda,$
where $\lambda$ is a finite real number. It follows from (7.3) that
$\lambda^{2}=2\lambda+t,$
whose zeros are $1\pm\sqrt{t+1}$. Since $r_{n}(1)>0$ for all $n\geq 2$, we have
(7.4) $\lambda=\lim_{n\rightarrow+\infty}\frac{f_{n+1}(1)}{f_{n}(1)}=1+\sqrt{t+1}.$
We proceed to compute $\mu_{n}$. By setting $x=1$ in the recursions (2.2), we have
$0\cdot f_{n+1}(1)$ $\,=-2n(t+1)f_{n}(1)+4(t+1)f_{n}^{\prime}(1),$
and hence,
(7.5) $\mu_{n}=\frac{f_{n}^{\prime}(1)}{f_{n}(1)}=\frac{2n(t+1)}{4(t+1)}=\frac{n}{2}.$
Letting $x=1$ in (2.3), we get
$4(n-1)(t+1)f_{n+1}(1)$ $\,=-4n(t+1)(nt+n-2t-3)f_{n}(1)+(4t+4)^{2}f_{n}^{\prime\prime}(1),$
Thus,
(7.6) $\frac{f_{n}^{\prime\prime}(1)}{f_{n}(1)}=\frac{n-1}{4(t+1)}\frac{f_{n+1}(1)}{f_{n}(1)}+\frac{n(nt+n-2t-3)}{4(t+1)}.$
It remains to compute $\sigma_{n}^{2}$ and its limitation as $n$ goes to +$\infty$+ ∞. By (7.5) and (7.6), we derived that
$\sigma_{n}^{2}=$ $\,\frac{f_{n}^{\prime\prime}(1)}{f_{n}(1)}+\mu_{n}-\mu_{n}^{2}$
$=$ $\,\frac{n-1}{4(t+1)}\frac{f_{n+1}(1)}{f_{n}(1)}+\frac{n(nt+n-2t-3)}{4(t+1)}+\frac{n}{2}-\frac{n^{2}}{4}$
$=$ $\,\frac{n-1}{4(t+1)}\frac{f_{n+1}(1)}{f_{n}(1)}-\frac{n}{4(t+1)}.$
By (7.4),
$\lim_{n\rightarrow+\infty}\frac{\sigma_{n}^{2}}{n}=$ $\,\lim_{n\rightarrow+\infty}\frac{n-1}{4n(t+1)}\frac{f_{n+1}(1)}{f_{n}(1)}-\lim_{n\rightarrow+\infty}\frac{n}{4n(t+1)}$
$=$ $\,\frac{1}{4(t+1)}\big{(}1+\sqrt{t+1}\big{)}-\frac{1}{4(t+1)}$
$=$ $\,\frac{1}{4\sqrt{t+1}}.$
So, we have $\sigma_{n}^{2}\sim\frac{n}{4\sqrt{t+1}}.$ and hence for $t>0$, $\lim_{n\rightarrow+\infty}\sigma_{n}^{2}=+\infty.$ This completes the proof. ∎
8. Laguerre inequality
In this section, we show the Laguerre inequalities arising from Speyer’s $g$-polynomial $g_{n,d}(t)$ and its generating function $h_{n}(x;t)$ for positive $t$. A polynomial $f(x)$ is said to satisfies the Laguerre inequality if
(8.1) $f^{\prime}(x)^{2}-f(x)f^{\prime\prime}(x)\geq 0.$
It was stated by Laguerre [34] that if $f(x)$ is a real-rooted polynomial, then the Laguerre inequality (8.1) holds for $f(x)$, see also [21].
Then we obtain the following result from Theorem 3.1.
Theorem 8.1.
Given $t>0$, the polynomial $f(x):=h_{n}(x;t)$ in $x$ satisfies the Laguerre inequality (8.1) for $n\geq 2$.
It was proved in [58, Theorem 3.1] that Speyer’s $g$-polynomial $g_{n,d}(x)$ has only real zeros. Hence, we get the following consequence.
Theorem 8.2.
The Speyer’s $g$-polynomial $g_{n,d}(x)$ satisfies the Laguerre inequality (8.1) for $n\geq 2$ and $11\leq d\leq n-$ 1.
Jensen [31] introduced a definition of generalized Laguerre inequality, that is, for $f(x)\in\mathcal{L}{\textrm{-}}\mathcal{P}$,
(8.2) $L_{r}(f(x))=\frac{1}{2}\sum_{k=0}^{2r}(-1)^{r+k}\binom{2r}{k}f^{(k)}(x)f^{(2r-k)}(x)\geq 0,$
where $f^{(k)}(x)$ is the $k$-th derivative of $f(x)$. Clearly, (8.2) reduces to the classical Laguerre inequality (8.1) for $r=1$. It is worth mentioning that Csordas and Vishnyakova [17] showed that if a function $f(x)$ satisfies (8.2) for all $r$ and all $x\in\mathbb{R}$, then $f(x)\in\mathcal{L}{\textrm{-}}\mathcal{P}$, which implies that the Laguerre inequality is a characterizing property of functions belonging to the Laguerre-Pólya class. For more properties on Laguerre inequality, see [53, 21, 56] and the references therein.
To conclude this section, we propose the following conjecture.
Conjecture 8.3.
Given $t>0$, the polynomial $h_{n}(x;t)$ in $x$ satisfies the generalized Laguerre inequality (8.2) for $r\geq 1$.
Acknowledgments
The author wishes to thank Arthur L. B. Yang for introducing him to this topic and for valuable discussions.